can u please draw and explain the v vs u graph for mirrors and lens
 

Grade:12

1 Answers

Vikas TU
6 years ago
The v versus u graph for mirror is always a hyperbola.
        As we know that,
       From the mirror formulae,
        1/u+1/v = 1/f
  •  (u+v)/uv = 1/f
  •  f(u+v) = uv
We know the equation of the type xy = a(x+y) asymptotic to x= a, y=a
Therefore the equation f(u+v) = uv is also asymptotic to u =f , v=f
And the graph comes as Hyperbola.